OpenStudy (anonymous):

rite an expression that represents the area of a circle with radius 3x^3. (For any circle with radius r, A = pi r^2 r times r

OpenStudy (anonymous):

*write

OpenStudy (anonymous):

Hi welcome to OPEnstudy! from the area formula substitute r=3x^3 then simplify like this: \(A= \pi r^2 =\pi (3x^3)^2\) then apply the power rule to simplify.. what will you get?

OpenStudy (anonymous):

\((x^m)^n=x^{mn}\)